# Locally Optimal Fixed-Budget Best Arm Identification in Two-Armed Gaussian Bandits with Unknown Variances

### 20 Dec 2023 v1

@@ 6.1 On the Asymptotic Optimality in Fixed-Budget BAI
6.2 Neyman Allocation with Unknown Variances
6.3 The AIPW, IPW, and Sample Average Estimators
A key component of our analysis is the AIPW estimator, which comprises an MDS and boasts minimum asymptotic variance. By using the properties of an MDS, we tackle the dependence among observations. The upper bound can also be applied to the IPW estimator, but in this case, the upper bound may not coincide with the lower bound. This discrepancy occurs because the AIPW estimator’s asymptotic variance is smaller than that of the IPW estimator. The minimum variance property of the AIPW estimator stems from the efficient influence function (Hahn, 1998; Tsiatis, 2007).
We conjecture that the asymptotic optimality of strategies employing the naive sample average estimator in the recommendation rule can be demonstrated, although we do not prove it in this study. This is because Hahn et al. (2011) shows that, using the CLT, the AIPW and sample average estimators have the same asymptotic distribution. However, due to the inability to utilize MDS properties and the presence of sample dependency, the analysis becomes challenging when we derive a corresponding result for a large deviation (exponential rate of the probability of misidentification). The findings of Hirano et al. (2003) and Hahn et al. (2011) may aid in resolving this issue.

### 17 Mar 2024 v2

@@ 7.2 Extension to BAI in Multi-Armed Bandit (MAB) Problems
8 Simulation Studies
This section provides simulation studies to investigate the empirical performance of the NA-AIPW strategy. For comparison, we also investigate the performances of the NA-IPW and NA-SA strategies defined in Section 6.3. Furthermore, we also conduct simulation studies of the “oracle” strategy with the known variances, denoted by Oracle, and the uniform strategy that draws an equal number of arms, denoted by Uniform. We recommend an arm with the highest sample average in the Oracle and Uniform strategies. [^2] The oracle strategy is the one proposed by Glynn & Juneja (2004) and Kaufmann et al. (2016). The Uniform strategy with the recommendation rule is referred to as the Uniform-Empirical Best Arm (EBA) strategy by Bubeck et al. (2011).
Throughout the experiment, we set arm $1$ as the best arm. We conduct experiments with the three settings.
In the first experiment, we set $\mu^{*}_{1}=1.00$ and choose $\mu^{*}_{2}$ from the set $\{0.80,0.85,0.90,0.95,0.99\}$. The variances $(\sigma_{1},\sigma_{2})$ are selected with a probability of $1/2$ from either $(1,v_{2})$ or $(v_{2},1)$, where $v_{2}$ is chosen from ${5,10,20,50}$. We continue the strategies until $T=10,000$ and report the empirical probability of misidentification at $T\in\{100,200,300,\dots,9,900,10,000\}$. [^3] This means we report the empirical probability of misidentification at $T\in\{100,200,300,\ldots,9,900,10,000\}$. We conduct $1,000$ independent trials for each choice of parameters and plot the results in Figures 1 and 2.
In the second experiment, the variances $(\sigma_{1},\sigma_{2})$ are selected with a probability of $1/2$ from either $(5,v_{2})$ or $(v_{2},5)$, where $v_{2}$ is chosen from ${5,10,20,50}$. The other settings are the same as the first experiment. The results are shown in Figures 3 and 4.
In the third experiment, we set $\mu^{*}_{1}=10.00$ and choose $\mu^{*}_{2}$ from the set $\{9.80,9.85,9.90,9.95,9.99\}$. The other settings are the same as the first experiment. The results are shown in Figures 5 and 6.
Our theoretical results imply that the probability of misidentifications of the NA-AIPW and Oracle strategies approach the same as $\Delta\to 0$. We can confirm the phenomenon. In the results, the Oracle strategy is a bit better than the NA-AIPW strategy when $\Delta$ is large. However, the gap approaches zero as $\Delta\to 0$. Note that when $\Delta$ is large, the convergence of the probability of misidentification is very fast, so the gap is still not so large even if $\Delta$ is large because both the probability of misidentifications of the NA-AIPW and Oracle strategies converge to zero very fast.
We can also find that the performance improvement of the NA-AIPW strategy from the Uniform strategy is large as the difference of variances is large. For example, in the second experiment, when $(\sigma_{1},\sigma_{2})=(5.5)$, there is no improvement by using the NA-AIPW strategy from the Uniform strategy because the NA allocation also leads us to draw each arm with equal ratio.
The difference between the NA-AIPW and NA-IPW strategies becomes large as the mean outcome of each arm becomes large. We can find that in the third setting, the NA-IPW strategy behaves badly since $\mu^{*}_{1}=10.00$ and $\mu^{*}_{2}$ is chosen from $\{9.80,9.85,9.90,9.95,9.99\}$, while $\mu^{*}_{1}=1.00$ and $\mu^{*}_{2}$ is chosen from $\{0.80,0.85,0.90,0.95,0.99\}$ in the first and second settings.
Figure 1: The results are under the first setting. We set $\mu^{*}_{1}=1.00$ and choose $\mu^{*}_{2}$ from the set $\{0.80,0.85,0.90,0.95,0.99\}$. The variances $(\sigma_{1},\sigma_{2})$ are selected with a probability of $1/2$ from either $(1,v_{2})$ or $(v_{2},1)$, where $v_{2}$ is chosen from ${5,10}$. We conduct $1,000$ independent trials and report the empirical probability of misidentification at $T\in\{100,200,300,\dots,9,900,10,000\}$.
Figure 2: The results are under the first setting. We set $\mu^{*}_{1}=1.00$ and choose $\mu^{*}_{2}$ from the set $\{0.80,0.85,0.90,0.95,0.99\}$. The variances $(\sigma_{1},\sigma_{2})$ are selected with a probability of $1/2$ from either $(1,v_{2})$ or $(v_{2},1)$, where $v_{2}$ is chosen from ${20,50}$. We conduct $1,000$ independent trials and report the empirical probability of misidentification at $T\in\{100,200,300,\dots,9,900,10,000\}$.